Assume that Selling Division and Buying Division are both owned by Overall Corporation. Selling Division sells a product that is used by Buying Division and outside customers. Selling Division has 22,000 units of excess capacity. Selling Division currently sells the product for $80 per unit and Buying Division currently buys 22,000 units of the product from an outside source for $80 per unit. Variable costs of the product are $16, of which $4 is the cost of selling the product to an outside customer.

Using Selling price less avoidable costs as the minimum price, fill in the following formula for the desired transfer price: $____< transfer price < $___________.

Using Variable costs as the minimum price, fill in the following formula for the desired transfer price: $________ < transfer price < $__________.

Assume there are no avoidable costs with an internal sale (variable costs equal $16) and that Buying Division buys 22,000 units from Selling Division. Complete the table for each transfer price:

Transfer Price Transfer Price
$75 $23
Increase in net income of Selling Division $ $
Increase in net income of Buying Division $ $
Increase in net income of Overall Corporation $ $
Answer #1

1. Using Selling price less avoidable costs as the minimum price, fill in the following formula for the desired transfer price: $ 76 < transfer price < $ 80.

Explaination: Variable selling price is avoidable cost. Hence selling price minus avoidable cost is 80-4 =76. Secondly maximum buying division is willing to pay is the price at which it is buying the product from market ie 80.

2. Using Variable costs as the minimum price, fill in the following formula for the desired transfer price: $ 16 < transfer price < $ 80.

Explaination: Since the variable selling cost are avoidable, we have other variable cost of 16 $. This will be the minimum price. Anything lesser than this will render loss to selling division. Hence it has to recover the costs incurred by it. The maximum is explained in the above.

3.

Transfer Price Transfer Price
$75 $23
Increase in net income of Selling Division $ (75 -16)*22000 = 1,298,000 $ (23-16)*22000 = 154000
Increase in net income of Buying Division $ (80-75)*22000 = 110,000 $ (80-23)*22000 = $ 1254,000
Increase in net income of Overall Corporation $ 1298,000+110,000 =1408000 $ 154000+1254000 = 1408000

Thus, no matter what the transfer price is. Overall profit to the business remains the same.